OpenStudy (anonymous):

Please help! Algebra 2 Find all zeros of the polynomial function 3. f(X)= X^4- 5X^2-36

OpenStudy (campbell_st):

well it's an equation that is reducible to a quadratric... so think about \[x^4 = (x^2)^2\] so you get \[f(x) = (x^2)^2 -5(x^2) - 36\] this can be factorised to \[f(x) = (x^2 - 9)(x^2 + 4)\] now you can find the solutions